Find all points where f fails to be differentiable. Justify your answer.

(a) f(x) = |3x − 2| (b) f(x) = |x^(2) −2|


1
Expert's answer
2021-10-20T16:33:19-0400

(a) "f(x)=3x-2,\;\;x>\frac{2}{3}."

"f(x)=2-3x,\;\;x<\frac{2}{3}."

f(x) is not differentiable at the point "(\frac{2}{3},0)".


(b) "f(x)=x^2-2, \;\;x<-\sqrt{2} \;or\; x>\sqrt{2}".

"f(x)=2-x^2 ,\;\;-\sqrt{2}<x<\sqrt{2}" .

f(x) is not differentiable at the points "(-\sqrt{2},0)" and "(\sqrt{2},0)."